one hundred pieces of a certain yarn were tested for strength, the mean value of the test results being 60N and their standard deviation 3N. calculate 95% confidence limits for the true mean strength of the yarn. if it was required to estimate the true mean correct to 1/2N . how many tests should be testes should be performed?


1
Expert's answer
2021-05-17T02:28:02-0400

Given, mean "\mu=60N,\sigma=3N"

"\alpha=0.05"


"\dfrac{\alpha}{2}=0.025, Z_{\frac{\alpha}{2}}=z_{0.025}=1.96"


95% confidence limit for true mean strength is given by-

"= \mu\pm Z_{\frac{\alpha}{2}}\dfrac{\sigma}{\sqrt{n}}
\\


 =60\pm 1.96\times \dfrac{3}{\sqrt{100}}
\\
 =60\pm 0.588
\\
 =(59.412,60.588)"


If we required to estimate the true mean correct to "\dfrac{1}{2}N" , Then The number of test we would have to perform is two, The first is the t-test for independent sample and second test is p-value test.